OpenStudy (anonymous):

-2x+7y=-9 7x+14y=-7 is the answer (1,-1),(-1,-1),(1,1) or (-1,1)?

mathslover (mathslover):

Multiply the first equation by 2 : -2x + 7y = - 9 ( multiply 2 both sides) It will become -4x + 14y = -18 Now, subtract this equation from the second equation : 7x + 14y - (-4x + 14y) = -7 - (-18) 7x + 14y + 4x - 14y = -7 + 18 Solve it further by cancelling out 14y.

OpenStudy (anonymous):

so would the set be (-1,1)?

mathslover (mathslover):

7x + 4x = 11 11x = 11 x = 1 and -2x + 7y = -9 -2(1) + 7y = -9 -2 + 7y = -9 7y = -9 + 2 7y = -7 y = -1 so, it will be (1,-1)

mathslover (mathslover):

You can always verify your result. Like you said that will it be (-1,1) ? then verify it.. you have 2 equations with you: -2x+7y=-9 7x+14y=-7 put x = -1 and y = 1 i) -2(-1) + 7(1) = -9 ; 2 + 7 = -9 ; 9 = -9 9 is not equal to -9.. as we already know. So, (-1,1) will not be our answer.

OpenStudy (anonymous):

thanks

mathslover (mathslover):

You're welcome :)